OpenStudy (b.r):

Is -7 + 9 = -9 + 7 true, flase, or open?

OpenStudy (zyi6):

True, I am guessing.

OpenStudy (shaik0124):

when u solve the answer u get is 2= -2 are this equal?

OpenStudy (anonymous):

its false -7+9= 2, -9+7=-2

OpenStudy (zyi6):

lol, read it wrong, ^ mrquesst is right.